3. Given the following information for Huntington Power Co., find the WACC. Assume the tax rate is 21%. a) Firm has 4,000 bonds with par value of $1,000, which are currently trading at $1,030 and has a maturity of 20 years. This bond makes semi-annual coupon payments of 7 percent.. b) Firm has 90,000 common shares outstanding, which are currently trading at $57 per share. Beta of the stock is 1.10. Market risk premium is 8 percent and risk-free rate is 6 percent. c) Firm has 13,000 preferred stock that pays $6 dividend. The current price of preferred stock is $104.
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=57*90000
=5130000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*4000*1.03
=4120000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=104*13000
=1352000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=5130000+4120000+1352000
=10602000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 5130000/10602000
W(E)=0.4839
Weight of debt = MV of Bond/MV of firm
Weight of debt = 4120000/10602000
W(D)=0.3886
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 1352000/10602000
W(PE)=0.1275
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 6 + 1.1 * (8)
Cost of equity% = 14.8
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =20x2
1030 =∑ [(7*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^20x2
                   k=1
YTM = 6.7249989046
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6.7249989046*(1-0.21)
= 5.312749134634
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 6/(104)*100
=5.77
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=5.31*0.3886+14.8*0.4839+5.77*0.1275
WACC =9.96%